# From Hyperons to Hypernuclei: A New Route to Unravel Proton Spin Polarization

### v3

@@ Abstract
Quark-gluon plasma, spin polarization, hypernuclei production, spin-orbit interaction, coalescence
@@
Figure 2 (a) shows the energy dependence of the global polarization of mid-rapidity ($|y|<1$) $\Lambda$ hyperons along the direction of total angular momentum at hadronization (blue band) and kinetic freeze-out stages (orange band) in collisions at impact parameter $b=8$ fm, corresponding to 20-50% centrality. For comparison, experimental data from the STAR Collaboration are represented by filled symbols for $\Lambda$ hyperons [2, 24] with combined statistical and systematic uncertainties. The global spin polarization of $\Lambda$ hyperons decreases with increasing beam energies, in good quantitative agreement with the STAR data as well as other theoretical calculations summarized in Refs. [28, 6, 29]. Note that the experimental data points are rescaled using the updated decay parameter of $\Lambda$ hyperon $\alpha_{\Lambda}=-\alpha_{\bar{\Lambda}}=0.732\pm 0.014$ [30].
Our results show that the $\Lambda$ polarization at kinetic freeze-out is reduced by approximately 20% relative to its value at hadronization. This suppression arises from combined effects of resonance feed-down and spin-flip scatterings during the hadronic phase. Specifically, the feed-down of primary $\Sigma^{*}$ resonances tends to enhance the $\Lambda$ polarization [31], while processes such as $\pi\Lambda\rightarrow\Sigma^{*}(1385)\rightarrow\pi\Lambda$ scattering and $\Sigma^{0}\rightarrow\gamma\Lambda$ decays lead to a reduction. The obtained magnitude of suppression from hadronic re-scatterings and decays is consistent with recent studies using rate equations [32, 33].
@@
Figure 3 (a) compares the model-calculated proton spin polarization (blue band) with the value of $(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}+\mathcal{P}_{\Lambda})/4$ (orange band), obtained from the calculated polarizations of $^{3}_{\Lambda}\text{H}$ and $\Lambda$ hyperons in Au+Au collisions. The two results are in excellent agreement, and their ratio, shown in the lower panel, remains consistent with unity within uncertainties. This confirms the validity of Eq. (1) as a reliable method to infer nucleon polarization from $^{3}_{\Lambda}\text{H}$ and $\Lambda$ measurements.
@@
Conclusions The spin information of protons and neutrons-the dominant baryonic constituents in nuclear collisions-is largely inaccessible in nuclear collisions due to their stabilities. To enable such measurements, we have proposed a new route to unravel the proton spin polarization through independent measurements of the polarizations of $\Lambda$ hyperons and $^{3}_{\Lambda}\text{H}$. To validate this method, we perform transport model calculations incorporating spin-dependent scatterings and resonance feed-down effects during the hadronic phase of Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV. We find that the proton polarization is related to the measurable $\Lambda$ and $^{3}_{\Lambda}\mathrm{H}$ polarizations through a simple and robust linear relation given by $\mathcal{P}_{p}\approx\left(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}+\mathcal{P}_{\Lambda}\right)/4$.

### v4

@@ Abstract
Quark-gluon plasma, spin polarization, hypernuclei production, spin-orbit coupling, hadronization
@@
Figure 2 (a) shows the energy dependence of the global polarization of mid-rapidity ($|y|<1$) $\Lambda$ hyperons along the direction of total angular momentum at hadronization (blue band) and kinetic freeze-out stages (orange band) in collisions at impact parameter $b=8$ fm, corresponding to 20-50% centrality. For comparison, experimental data from the STAR Collaboration are represented by filled symbols for $\Lambda$ hyperons [2, 24] with combined statistical and systematic uncertainties. The global spin polarization of $\Lambda$ hyperons decreases with increasing beam energies, in good quantitative agreement with the STAR data as well as other theoretical calculations summarized in Refs. [28, 6, 29]. Note that the experimental data points are rescaled using the updated decay parameter of $\Lambda$ hyperon $\alpha_{\Lambda}=0.732\pm 0.014$ [30].
Our results show that the $\Lambda$ polarization at kinetic freeze-out is reduced by approximately 20% relative to its value at hadronization. This suppression arises from combined effects of resonance feed-down and spin-flip scatterings during the hadronic phase. Specifically, the feed-down of primary $\Sigma^{*}$ resonances tends to enhance the $\Lambda$ polarization [31], while processes such as $\pi\Lambda\rightarrow\Sigma^{*}(1385)\rightarrow\pi\Lambda$ scatterings and $\Sigma^{0}\rightarrow\gamma\Lambda$ decays lead to a reduction. The obtained magnitude of suppression from hadronic re-scatterings and decays is consistent with recent studies using rate equations [32, 33].
@@
Figure 3 (a) compares the model-calculated proton spin polarization (blue band) with the value of $(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}+\mathcal{P}_{\Lambda})/4$ (orange band), obtained from the calculated polarizations of $^{3}_{\Lambda}\text{H}$ and $\Lambda$ hyperons in Au+Au collisions. The two results are in excellent agreement, and their ratio ($(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}+\mathcal{P}_{\Lambda})/4\mathcal{P}_{p}$), shown in the lower panel, remains consistent with unity within uncertainties. This confirms the validity of Eq. (1) as a reliable method to infer nucleon polarization from $^{3}_{\Lambda}\text{H}$ and $\Lambda$ measurements.
@@
Conclusions The spin information of protons and neutrons-the dominant baryonic constituents in nuclear collisions-is largely inaccessible due to their stabilities. To enable such measurements, we have proposed a new route to unravel the proton spin polarization through independent measurements of the polarizations of $\Lambda$ hyperons and $^{3}_{\Lambda}\text{H}$. To validate this method, we perform transport model calculations incorporating spin-dependent scatterings and resonance feed-down effects during the hadronic phase of Au+Au collisions at $\sqrt{s_{NN}}=7.7$–200 GeV. We find that the proton polarization is related to the measurable $\Lambda$ and $^{3}_{\Lambda}\mathrm{H}$ polarizations through a simple and robust linear relation given by $\mathcal{P}_{p}\approx\left(3\mathcal{P}_{\text{${}^{3}_{\Lambda}\text{H}$}}+\mathcal{P}_{\Lambda}\right)/4$.